OpenStudy (anonymous):

Find |u+v| given that Theta (θ) is the angle between |u| and |v| a.|u|=66, |v|=22, angle θ=44°

OpenStudy (anonymous):

try finding \[|u+v|^{2}=?\] and then put a square on it

OpenStudy (anonymous):

how does the angle come into play

OpenStudy (anonymous):

it must involve some multiplication with sine or cosine

OpenStudy (anonymous):

\[|u+v|^{2}=u^{2}+v^{2}+2uv(*)=|u|^{2}+|v|^{2}+|u||v|\cosθ=...\] u*v=|u||v|cosθ (vector product)

OpenStudy (anonymous):

and then plug the result u got at a square root to find the |u+v|

OpenStudy (anonymous):

u got it?

OpenStudy (anonymous):

im not getting the right answer

OpenStudy (anonymous):

i dont understand the equation you typed out. i cant fathom it without numbers in it. its hard for me to understand things when steps are in a string

OpenStudy (anonymous):

its easier to see equivalency when things are stacked when simplified or expanded

OpenStudy (anonymous):

\[|u+v|^{2}=(u+v)^{2}=u^{2}+2uv+v^{2}=|u|^2+2uv(i)+|v|^{2} \] for any vector: \[a^{2}=|a|^{2}\] to get rid of (i) we use the vector product